Q4D.1 (Requires completion of chapter R9.) If we shine X-rays (which are very high-frequency electromagnetic waves) at a target, both the wave and photon models of electromagnetic radiation predict that the X-rays will be scattered in all directions by the target. The wave model predicts that the scattered X-rays will have the same wave- length as that of the incoming X-rays (since the incoming waves will wiggle the electron at the same frequency, so the scattered waves emitted by the electron in all direc- tions must have the same frequency as the bobbing elec- n that emits them). The photon model, on the other hand, predicts that the wavelength of the scattered X-rays will depend on the angle in a specific way. Our purpose here is to derive this angle dependence Lets define our coordinate system so that the initial X-ray photons move in the +x direction. Let the initial energy of the X-ray photons be Eo and the mass of an elec- tron be m. Imagine that a photon strikes the electron and scatters at an angle 0, while the electron scatters at an angle P, as shown below. BEFORE AFTER: photon,y energy Eo o photon energy E electron at rest scattered electron

media%2Fc9c%2Fc9ca20f1-9ed0-4746-bf1c-46
Answer #1

As per the equation given in the equation of matrices, m+E, - E+ m2 + p.1) Photon incident is given as E = ln , and photon sc

b). According to the condition given in part b of the question, E = E cose +|p.cos 0 Esine+p.sino p=p+ p. P. = p-p squaring

bv + m,c* = + vcp2 +m c* P+E = pc+VE +p?c? On Squaring both sides,we get E;+c(p-p)*+2CE (P-p) = E + pc? p = p + po-2 pp